OpenStudy (anonymous):

line that has a slope os 2/3 and y intecept of -4

OpenStudy (anonymous):

y=2x/3-4

OpenStudy (anonymous):

y=2/3x-4

OpenStudy (amistre64):

\[y = \frac{2}{3}x-4\]

OpenStudy (anonymous):

y=mx+c where m is the slope and c is the y-intercept is the genral form of line equation.

OpenStudy (anonymous):

okay i'm lostttt!

OpenStudy (anonymous):

y=2/3x-4

OpenStudy (anonymous):

my options are y=2/3x+4,y=-4x+2/3,y=2/3x-4,y=-2/3x-4

OpenStudy (anonymous):

It is y=2/3x-4

OpenStudy (anonymous):

bubblegum97 YOU WAS RIGHT!!!!

OpenStudy (anonymous):

Lol, I was ! lmao YAY ...Were learning this stuff in my school, Im in 7th grade. Sounds good to hear that my answer is correct :)